# RM-R1: Reward Modeling as Reasoning

### v1

@@ 2 Related Work
3 RM-R1
Figure 2: Training pipeline of RM-R1. Starting from a generative reward model (GenRM), RM-R1 training involves two stages: Distillation and Reinforcement Learning (RL). In the Distillation stage, we use high-quality synthesized data to bootstrap RM-R1’s reasoning ability. In the RL stage, RM-R1’s reasoning ability for reward modeling is further strengthened. After distillation, a GenRM evolves into a ReasRM. RM-R1 further differentiates itself by being RL finetuned on preference data.
Figure 2 illustrates the overall training pipeline of RM-R1. Starting from any off-the-shelf instruction-tuned model (e.g., Qwen-2.5-14b-instruct), we synthesize high-quality reasoning traces and distill RM-R1 on the synthesized reasoning traces. After distillation, RM-R1 unlocks the basic reasoning ability for reward modeling. However, distilled models often still suffer from overfitting to certain patterns in the distillation data, thus, the model’s ability to generalize may be limited. To further improve RM-R1’s performance, we use RL to continually train and derive the full model of RM-R1. In this section, we begin by defining the reward modeling task (Section 3.1), then briefly describe the distillation process (Section 3.2). Later, we introduce in detail how we conduct RL training, especially our customized rollout strategy and reward definition (Section 3.3).
3.1 Task Definition
Motivation: Collecting explicit scalar rewards for LLM outputs from human annotators is often expensive and impractical. It is challenging to assign consistent numerical scores due to factors like subjective biases and variable interpretation of scales. Conversely, obtaining pairwise preference judgments—simply indicating which of two responses is better—is significantly more intuitive and reliable for annotators. Such preference data is easier to acquire at scale and inherently aligns with human decision-making processes, resulting in more consistent and actionable training signals.
Motivated by this practicality, recent approaches advocate training reward models using pairwise preference datasets. Among them, training GenRMs in an LLM-as-a-judge fashion has emerged as particularly promising [26, 39, 41]. Unlike traditional classifiers or regressors that output single scalar values, generative reward models produce comprehensive, textual judgments, including rationales and explicit reasoning about their preferences. This approach leverages LLM’s natural ability for nuanced evaluation and interpretability, effectively capturing complex, multi-dimensional criteria beyond simplistic scalar rankings.
Formal Definition: Given a preference dataset:
$\mathcal{D}=\{(x^{(i)},y_{a}^{(i)},y_{b}^{(i)},l^{(i)})\}_{i=1}^{N},$
where $x$ is a prompt, $y_{a}$ and $y_{b}$ are two different responses for $x$. $l\in\{a,b\}$ is the ground truth label that indicates the preferred response. We define the generative reward modeling task as follows:
Let $r_{\theta}$ denote a generative reward model parameterized by $\theta$. For each data sample $(x^{(i)},y_{a}^{(i)},y_{b}^{(i)},l^{(i)})$, $r_{\theta}$ generates a textual judgment $j$ consisting of tokens $j=(j_{1},j_{2},\dots,j_{T})$, modeled by:
Note that $j$ contains $r_{\theta}$’s prediction of the preferred response $\hat{l}\in j$. The goal of reward modeling is :
3.2 Distillation of Reasoning Trace
For an instruct model, e.g., Qwen-2.5-14b-instruct [38], it is quite intuitive to turn it into a GenRM simply by prompting. However, without fine-tuning on reward modeling reasoning traces, these models may struggle to conduct consistent judgments. To bootstrap an instruct model’s reasoning potential, we start with training an instruct model with long reasoning traces synthesized for reward modeling. Specifically, we subsample $M$ data samples from $\mathcal{D}$ and denote it as $\mathcal{D}_{sub}$. Given a data sample $(x^{(i)},y_{a}^{(i)},y_{b}^{(i)},l^{(i)})\in\mathcal{D}_{sub}$, we ask an “oracle” model like o3 or claude-3-7-sonnet to generate its structured reasoning trace $r^{(i)}$ justifying why $y_{l}^{(i)}$ is chosen as the correct answer of $x^{(i)}$. We then construct the reasoning trace ground truth:
$y_{\mathrm{trace}}^{(i)}=r^{(i)}\oplus y_{l}^{(i)},$ (2)
$\mathcal{D}_{\mathrm{distill}}=\{(x^{(i)},y_{\mathrm{trace}}^{(i)})\}_{i=1}^{M},$ (3)
$\mathcal{L}_{\mathrm{distill}}(\theta)=-\sum_{(x,y)\in\mathcal{D}_{\mathrm{distill}}}\sum_{t=1}^{T}\log r_{\theta}\left(y_{t}\mid x,y_{<t}\right)$ (4)
where $y_{<t}=(y_{1},y_{2},...,y_{t-1})$ denotes the sequence of tokens preceding position $t$.
3.3 RL Training
Although distilling is a proper way to turn a general generative model into a GenRM, it often suffers from overfitting to certain patterns and constrains the model’s ability to generalize its reasoning abilities for critical thinking [32] , which is essential for reward modeling. To address this, we propose to integrate RL as a more powerful learning paradigm to enhance the model’s ability to conduct reward-based reasoning. Training a policy model using RL has been widely seen in the preference optimization phase of LLM post-training [25], and it is quite natural to extend this paradigm for training a ReasRM. To be specific, we directly treat our reward model $r_{\theta}(j\mid x,y_{a},y_{b})$ as if it is a policy model:
where $r_{\mathrm{ref}}$ is the reference reward model. In practice, we use the checkpoint before training as $r_{\mathrm{ref}}$, and that means $r_{\mathrm{ref}}$ could be an off-the-shelf LLM or the LLM obtained after the distillation step in Section 3.2. $\mathcal{R}(x,j)$ is the reward function, and $\mathbb{D}_{\mathrm{KL}}$ is KL-divergence. The $x$ denotes input prompts drawn from the preference data $\mathcal{D}$. The $j$ indicates the text generated by the reward model, which includes the reasoning trace.
3.3.1 Chain-of-Rubrics Rollout
To facilitate the distilled models to proactively generate critical reasoning traces, we design a system prompt as shown in Figure 3 during rollout. Intuitively, reward modeling for general domain (chat, safety, etc.) and reasoning domain (math, code, etc.) should focus on different angles. For example, for the chat domain, we may care more about some aspects that can be expressed in textual rubrics (e.g., be polite), yet for the reasoning domain, we usually care more about logical coherence and answer correctness. Based on this intuition, we instruct ${r_{\theta}}$ to classify each preference data sample $\{(x,y_{c},y_{r})\}$ into one of the two <type>: Chat or Reasoning. For each <type>, we ask ${r_{\theta}}$ to practice different reasoning step by step: For reasoning tasks, we ask ${r_{\theta}}$ to solve $x$ on its own. During the <eval> phase, $r_{\theta}$ compares $y_{c}$ and $y_{r}$ conditioned on its own </solution> and selects an <answer>. Regarding the Chat type, we instead ask ${r_{\theta}}$ to think about and justify the <rubric> for grading the chat quality (including safety). Since Chat scenarios do not have a verifiable reward, we instruct the model to select the <answer> based on its own <rubric>.
@@ 3.3.1 Chain-of-Rubrics Rollout
2. Assign weights to each rubric item based on their relative importance. 3. Inside <rubric>, include a <justify>…</justify> section explaining why you chose those rubric criteria and weights. 4. Compare both Chatbot responses according to the rubric. 5. Provide your evaluation inside <eval>…</eval> tags, using <quote_A>, <summary_A>, <quote_B>, and <summary_B> as described above. 6. End with your final judgment in the format: <answer>[[A]]</answer> or <answer>[[B]]</answer> Important Notes:
- Be objective and base your evaluation only on the content of the responses.
- Do not let response order, length, or Chatbot names affect your judgment. - Follow the response format strictly depending on the task type.
Figure 3: The system prompt used for RM-R1 rollout (for instruct models).
@@
3.3.2 Reward Design
Rule-based reward mechanisms have demonstrated strong empirical performance to facilitate reasoning since DeepSeek-R1 [11], and have been commonly employed ever since. In our training, we further simplify the reward formulation of DeepSeek-R1 and merely focus on correctness-based component, in line with prior works [30, 19]. We have also tried adding the format reward to the overall reward, but found that the task performance does not have a significant difference. The rationale behind only focusing on correctness is that the SFT models have already learned to follow instructions and format their responses properly. During experiments, we also observe that adding a format reward makes the advantage computation significantly slower, since regular expression matching grows linearly with the response token length.
Formally, our reward is defined as follows:
$\mathcal{R}(x,j|y_{a},y_{b})=\begin{cases}1,&\vbox{\hbox{\parbox{156.49014pt}{if $\hat{l}=l$}} \vspace{8pt}}\\ -1,&\text{otherwise.}\end{cases}$ (6)
3.3.3 Group Relative Policy Optimization (GRPO)
@@
$\begin{split}\mathcal{J}_{\text{GRPO}}(\theta)=\,&\mathbb{E}_{x\sim\mathcal{D},\,\{j_{i}\}_{i=1}^{G}\sim r_{\theta_{\text{old}}}(j\mid x)}\bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|j_{i}|}\sum_{t=1}^{|j_{i}|}\Big{\{}\min\Big{(}\frac{r_{\theta}(j_{i,t}\mid x,j_{i,<t})}{r_{\theta_{\text{old}}}(j_{i,t}\mid x,j_{i,<t})}\hat{A}_{i,t},\\[4.0pt] &\text{clip}\big{(}\frac{r_{\theta}(j_{i,t}\mid x,j_{i,<t})}{r_{\theta_{\text{old}}}(j_{i,t}\mid x,j_{i,<t})},1-\epsilon,1+\epsilon\big{)}\hat{A}_{i,t}\Big{)}-\beta\,\mathbb{D}_{\mathrm{KL}}\left[r_{\theta}(\cdot\mid x)\,\|\,\pi_{\text{ref}}(\cdot\mid x)\right]\Big{\}}\bigg{]}\end{split}$ (7)
where $\beta$ is a hyperparameter balancing the task specific loss and the KL-divergence. Specifically, $\hat{A}_{i}$ is computed using the rewards of a group of responses within each group: $\{r_{1},r_{2},\ldots,r_{G}\}$, and is given by the following equation:
@@
4 Experiments
4.1 Experimental Setup
4.1.1 Benchmarks
@@
RewardBench [18]: RewardBench is one of the first endeavors towards benchmarking reward models through prompt-chosen-rejected trios, covering four categories: chat, chat-hard, reasoning, safety, with 358. 456, 740. 1431 samples, respectively.
@@
4.2 Preference Datasets
We consider the following datasets for training (A summary is presented in Appendix B):
@@ 4.2 Preference Datasets
4.3 Baselines
@@
GenRMs. Generative reward models provide more nuanced feedback by generating textual judgments. Examples include Llama [9], Qwen [38], Claude [3], GPT-4 variants [1, 14], Gemini-1.5-pro [27] and Self-taught Evaluators [36]. GenRMs leverage generative capabilities of LLMs to produce explanations and rationales alongside numerical scores, enhancing interpretability and decision transparency.
ReasRMs. Reasoning-enhanced reward models explicitly incorporate reasoning processes into their judgments, often trained through critiques or chain-of-thought strategies. Notable examples are Critique-RM [39], DeepSeek-GRM [24] and our proposed RM-R1 models. These models excel in tasks demanding rigorous reasoning, safety evaluations, and nuanced preference judgments due to their grounding in structured critical thinking.
4.4 Implementation Details
Our training framework is based on verl [31] and OpenRLHF [13]. We use the training sets detailed in Section 4.3. For Instruct models, we use 8.7k data for distillation and 64k for RLVR. For Deepseek-Distilled models, we use the full data for RLVR. The full implementation details are provided in Appendix C.
4.5 Main Results
For the baselines, we try our best to reproduce the numbers if essential resources are open-sourced (e.g., model checkpoints, system prompts). Otherwise, we use the numbers reported in the corresponding tech report or benchmark leaderboard. For each benchmark, we cutoff and select the best-performing models in each category for brevity.
Table 1: Results of our proposed method and baselines on the RewardBench. Bold numbers indicate the best performance, Underlined numbers indicate the second best.
@@ 4.5 Main Results
RM-R1-Qwen-Instruct-32B 95.3 83.1 91.9 95.2 92.9
Table 1, Table 2, and Table 3 report the performance of RM-R1 and baseline models on RewardBench, RM-Bench, and RMB, respectively. Our key findings are summarized below:
RM-R1 Achieves State-of-the-Art Performance. We first compare our two strongest models—RM-R1-Qwen-Instruct-32B and RM-R1-DeepSeek-Distilled-Qwen-32B—against top-performing baselines. On Reward Bench, RM-R1-Qwen-Instruct-32B achieves state-of-the-art performance among all existing generative reward models, outperforming closed-source commercial systems such as GPT-4o, Gemini-1.5-Pro, and models up to 340B in size. On RM-Bench, RM-R1-DeepSeek-Distilled-Qwen-32B surpasses the previous state-of-the-art by up to 12.8%, marking a significant improvement. On RMB, RM-R1-Qwen-Instruct-32B attains near state-of-the-art performance with 73.0%, closely matching the best prior result of 73.8%. Notably, this model is trained on only 73K preference examples, demonstrating strong data efficiency.
Reasoning Training Demonstrates Effectiveness. Our reasoning-based training pipeline yields substantial gains. For example, RM-R1-Qwen-Instruct-14B consistently outperforms DeepSeek-GRM-27B, a reasoning model nearly twice its size. We provide a detailed analysis of the key training design choices that contribute to this success in Section 5.2, and further highlight the advantages of reasoning-based training over non-reasoning counterparts in Section 5.3.
RM-R1 Generalizes Across Domains. Our generalist reward models based on Qwen-2.5-Instruct exhibit strong performance across all domains, including chat, safety, and reasoning, with no observable domain bias. Meanwhile, models trained from DeepSeek-Distilled-Qwen benefit from extensive reasoning-focused supervision and achieve stronger performance in reasoning-centric benchmarks. Specifically, RM-R1-DeepSeek-Distilled-Qwen-32B sets a new state-of-the-art on RM-Bench with 91.8% in math and 74.1% in code, outperforming previous best models by substantial margins (66% in math and 59% in code). It also achieves the strongest reasoning performance among our released models on RewardBench. However, in terms of data efficiency, our Instruct-based models reach competitive reasoning performance and superior overall results using only 8.7K examples for distillation—compared to the 800K examples used in training DeepSeek-Distilled.
Additional Analysis. Across all benchmarks, we observe that larger models consistently achieve higher performance. This trend is not coincidental. In Section 5.1, we conduct a detailed analysis and show that reasoning reward models benefit more from increased model scale, reinforcing the effectiveness of scaling in this setting. We also conduct more fine-grained analysis of the training dynamics and case study of RM-R1 in Section 5.4 and 5.5.
@@
5 Analysis
In this section, we present a series of empirical studies and ablation experiments aimed at understanding the key ingredients for training effective reasoning reward models. Our analysis spans scaling effects, design decisions, reasoning ablations, training dynamics, and case studies.
5.1 Scaling Effects
We begin by investigating how model performance varies with scale, considering both model size and inference-time compute. While prior work has shown that larger models generally achieve lower pretraining loss with stronger capabilities [15], and recent studies highlight the benefits of increased inference budgets [15], reward modeling presents a more nuanced picture. In some cases—such as scalar-valued sequence models from Skywork [21]—the performance gap between the small model (e.g., 8b) and the large model (e.g., 27b) is minimal (often less than 1%), showing no clear advantage from scaling. In this subsection, we show that this trend does not hold for reasoning reward models, where scaling brings clear and substantial improvements.
5.1.1 Model Sizes
We now analyze the impact of model scale. Our study is based on the Qwen-2.5-Instruct model family at three sizes: 7B, 14B, and 32B. We evaluate performance improvements resulting from our training procedure described in Section 3, with results averaged across three key benchmarks: RewardBench, RM-Bench, and RMB.
For each model size, we compare the original and post-training performance. Figure 5(a) plots the absolute improvement with respect to the original model performance (in %) as a function of model size. Observing an approximately linear trend, we fit a linear regression model and extrapolate to hypothetical scales of 3B and 72B, shown using faint markers and dashed extensions. The results strongly support a scaling law for reasoning reward models (RM-R1): larger models not only result in an absolute better final performance but also consistently yield greater performance gains. This aligns with the intuition that our training effectively leverages the superior reasoning capabilities of larger models. Furthermore, these findings provide compelling evidence that reinforcement learning benefits from scale as well.
(a) Model Size
(b) Inference Compute
Figure 5: Scaling effect of RM-R1. We find that (a) larger models benefit more from reasoning training, and (b) longer reasoning chains result in better reward modeling performance.
5.1.2 Inference-time Computation
Next, we examine the effect of inference-time computation—how model performance varies with different compute budgets (measured in number of tokens allowed during inference). Since this is particularly relevant to reasoning-focused models, we fix our base model to DeepSeek-R1-Distill-Qwen-14B. We evaluate average performance across the three key benchmarks using a wide range of inference-time compute budgets: 512, 1024, 2048, 4096, and 8192 tokens.
To ensure a fair comparison, we match the training compute to the inference budget in each setting (i.e., we allow a maximum of $x$ tokens during training for a compute budget of $x$ at inference). All models are trained using GRPO with identical datasets and hyperparameter configurations. Figure 5(b) shows the relationship between compute budget and performance. We observe a clear improvement trend as the inference budget increases. This highlights the benefits of long reasoning chains in reward modeling.
[…]
5.2 Training Recipes
We next investigate the key ingredients underlying the successful training of RM-R1. Through a series of ablation studies, we examine our design choices to identify effective strategies for training high-quality reasoning reward models. We compare the following settings:
Cold Start RL. Prior work on reasoning model training typically adopts a straightforward pipeline based on reinforcement learning with verifiable rewards [5, 24]. This approach generally involves pure RL, with rule-based rewards centered on answer correctness and format compliance. Such strategies have achieved notable success in advanced mathematical problem solving.
@@ 5.2 Training Recipes
$\mathcal{R}_{\text{format}}=\begin{cases}1&\text{ if format matches}\\ 0&\text{ otherwise}\end{cases}\quad\quad\mathcal{R}_{\text{answer}}=\begin{cases}1&\text{ if answer matches}\\ 0&\text{ otherwise}\end{cases}$ (9)
@@
The total reward is the sum $\mathcal{R}=\mathcal{R}_{\text{answer}}+\mathcal{R}_{\text{format}}$. We use the prompt template shown in Figure 7, a simplified version of Figure 3. In this setup, all input prompts are treated uniformly—that is, chat and reasoning tasks are not distinguished.
Cold Start RL + Query Categorization (QC). This setting largely follows the previous one, with a key modification: prompting the LM to first categorize the task into reasoning or chat, and then apply different strategies for handling those tasks. Intuitively, reinforcement learning alone can effectively explore reasoning tasks, a domain where it has already achieved considerable success. Here, we incorporate the system prompt shown in Figure 3, which explicitly distinguishes between chat and reasoning tasks.
Distilled + RL + QC (RM-R1). Building on the previous setup, we introduce an additional distillation stage from stronger teacher models as a warm start before RL training. The motivation is that with RL alone, weaker models (especially at smaller scales) often fail to explore high-quality rubrics and convincing reasoning chains for chat tasks throughout the RL training process. Distilling strong reasoning traces on a small subset of data can effectively mitigate this limitation.
This two-stage process yields a high-quality distillation set. We deliberately choose the order—first Claude, then O3—based on qualitative observations: Claude excels at solving easier tasks and maintaining attention to safety considerations, whereas O3 performs better on harder helpfulness tasks but tends to overemphasize helpfulness at the expense of safety. We select approximately 12% of the training data (slightly fewer than 9K examples) for distillation. This is then followed by RL training.
Table 4: Ablation Study of the Designs of the Qwen-2.5-Instruct-32B on Reward Bench.
Method Chat Chat Hard Safety Reasoning Reward Bench
Instruct (Original) 95.8 74.3 86.8 86.3 85.8
Instruct + Cold Start RL 93.8 78.1 90.3 90.6 88.2
Instruct + Cold Start RL + QC 95.0 78.1 88.9 95.8 89.8
Instruct + Distilled + RL + QC 95.3 83.1 91.9 95.2 92.9
  • Query Categorization significantly boosts reasoning performance. Incorporating explicit query categorization into the prompt notably improves reasoning performance, suggesting that clearer task guidance benefits learning.
[…]
5.3 Effectiveness of Reasoning Training
We now analyze the impact of reasoning-based training. Throughout this subsection, we use SFTto refer to supervised fine-tuning on final answers only, and distilled to refer to supervised fine-tuning on reasoning chains generated by a stronger model. Prior work has shown that SFT alone can yield substantial performance gains [21]. Here, we demonstrate that reasoning-based training can outperform such answer-only approaches. We consider the following experimental settings:
Table 5: Comparison of reasoning-based training versus SFT, measured across the three key benchmarks. Methods marked with * involve reasoning training. We can find that reasoning training consistently results in a better-performing reward model.
Method RewardBench RM-Bench RMB Average
Full Data
Instruct + SFT 90.9 75.4 65.9 77.4
Instruct + Distilled + SFT 91.2 76.7 65.4 77.8
Instruct + RM-R1 * 92.9 79.1 73.0 81.7
9k Data
Instruct + SFT 88.8 74.8 66.9 76.6
Instruct + Distilled* 89.0 76.3 72.0 79.2
In summary, both (+ RM-R1) and (+ Distilled) represent reasoning-based approaches, while the remaining methods are purely non-reasoning-based approaches. In Table 5, we report the results measured across the three benchmarks. The findings clearly demonstrate that reasoning training significantly benefits reward model performance. Under fair comparisons (i.e., training on exactly the same amount of data), reasoning-based models consistently outperform their SFT-only counterparts. In particular, even high-quality distillation alone, applied to a small subset of the data, provides notable gains, highlighting the value of incorporating structured intermediate reasoning.
[…]
5.4 Training Dynamics
(a) Cold Start RL
(b) Warm Start RL
5.5 Case Study
Question: symptoms of sickle-cell disease Chatbot A: 1. Painful red or yellow skin lesions; 2. Frequent infections; 3. Fatigue and weakness; 4. Anemia; 5. Slower growth and development; 6. Jaundice; 7. Swelling of hands and feet; 8. Vision loss; 9. Tiredness and weakness after physical activity; 10. Breathlessness; 11. Fever, chills, and night sweats. Chatbot B Sickle cell disease is a genetic disorder that affects the shape of red blood cells. It is most common in African Americans, but can also affect people from other ethnic groups. The symptoms of sickle cell disease can vary widely from person to person, but some common symptoms include: 1. Fatigue: People with sickle cell disease often experience fatigue and weakness, even when they have not been physically active.; 2. Pain: Sickle cell disease can cause pain in the body, particularly in the arms, legs, back, and abdomen. Pain can be mild or severe and can last for days or weeks.; 3. Fever: People with sickle cell disease may be prone to infections and may experience frequent fevers.; 4. Anemia: Sickle cell disease can cause anemia, a condition in which the body does not have enough red blood cells to carry oxygen throughout the body.; 5. Delayed growth and development: Children with sickle cell disease may be smaller than their peers due to delayed growth and development.; 6. Infections: People with sickle cell disease are at increased risk for infections, particularly in the bloodstream and urinary tract.; 7. Pigmentation: Sickle cell disease can cause dark areas of pigmentation in the body, particularly on the hands and feet. It’s important to note that not everyone with sickle cell disease will experience all of these symptoms, and some people may not experience any symptoms at all. Therefore, it’s important to consult with a healthcare provider if you suspect you or someone you know may have sickle cell disease. Cold-start RL evaluation <type>Chat</type> <rubric> Relevance 40 % – Comprehensiveness 30 % – Clarity/organisation 30 % </rubric> <eval> Chatbot A provided a list of symptoms that are directly relevant to sickle cell disease. The response is comprehensive and includes a wide range of symptoms commonly associated with the condition. The symptoms are listed in a clear and organized manner, making it easy for the client to understand the information. Chatbot B’s response is less organized than Chatbot A’s response, and the explanation of sickle cell disease may be unnecessary for the client’s question.</eval> <answer>[[A]]</answer> RM-R1 evaluation <rubric> 1 Accuracy 40 % · 2 Comprehensiveness 30 % · 3 Clarity 20 % · 4 Helpfulness 10 % <justify> Accuracy is weighted highest because providing correct information about medical conditions is crucial. Comprehensiveness is important because …</justify> </rubric> <eval> Accuracy of information: Chatbot A provides a list of symptoms without explanation, but some listed items (e.g. “painful red or yellow skin lesions,” “vision loss”) are not typical of sickle-cell disease. Chatbot B offers more accurate information, explaining each symptom’s nature and severity. … Comprehensiveness:Clarity and organization:Helpfulness:</eval> <answer>[[B]]</answer>
Table 6: Case Study of RM-R1 (succinct version). Complete model responses are shown in Table 7 and Table 8. Cold-Start RL evaluation tends to focus on superficial features (e.g., number of suggestions, relevance), whereas RM-R1 produces high-quality rubrics (e.g., accuracy), faithfully follows them, and evaluates responses based on their actual content.
To gain deeper insights into RM-R1, we conduct a case study using Qwen-14B-Instruct, comparing its behavior with a cold-start RL model. The results, summarized in Table 6, reveal the following key observations:
Transparent Judging Process.
High-Quality, Question-Dependent Rubrics.
Faithful Adherence to Rubrics and Content-Based Judgement.
6 Conclusion and Future Work
In this paper, we revisited reward modeling through the lens of reasoning. We introduced RM-R1, a family of ReasRMs that effectively generate explicit chains of rubrics and rationales, and scale with both model size and inference compute. Across three public benchmarks, RM-R1 matched or surpassed commercial and open-source RMS while producing more interpretable judgments. Ablation investigations reveal that (1) task-type classification, (2) bootstrapping from high-quality reasoning traces, and (3) RL fine-tuning are all indispensable. Qualitative analyses further showed that RM-R1 learns to prioritize high-impact rubrics, faithfully follow its own criteria and justify coherently.
Future work includes automatic rubric induction, where an explicit rubric library can be reused and composed, so that rollout length could be reduced; Another direction is active preference collection, where ReasRMs use active learning to query human preference only when the current rubric set is insufficient for a new preference sample. Finally, it would be natural to extend our study to multimodal/agentic reward modeling scenarios.
Appendix A Supplementary Information for Section 5

### v2

@@ 5 Related Work
6 Conclusion and Future Work
In this paper, we revisited reward modeling through the lens of reasoning. We introduced RM-R1, a family of ReasRMs that effectively generate explicit chains of rubrics and rationales, and scale with both model size and inference compute. Across three public benchmarks, RM-R1 matched or surpassed commercial and open-source RMs while producing more interpretable judgments. Ablation investigations reveal that (1) task-type categorization, (2) bootstrapping from high-quality reasoning traces, and (3) RL fine-tuning are all indispensable. Qualitative analyses further showed that RM-R1 learns to prioritize high-impact rubrics, faithfully follow its own criteria and justify coherently. Future work includes active preference collection, where ReasRMs use active learning to query human preference only when the current rubric set is insufficient for a new preference sample. Finally, it would be natural to extend our study to multimodal/agentic reward modeling scenarios.
Acknowledgments and Disclosure of Funding
The research reported in this paper/presentation was performed in connection with the U.S. Army Contracting Command - Aberdeen Proving Ground (ACC-APG) and the Defense Advanced Research Projects Agency (DARPA) under contract number W912CG- 24-C-0011. The views and conclusions in this paper/presentation are those of the authors and should not be interpreted as presenting the official policies or position, either expressed or implied, of ACC-APG, DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation hereon.
Appendix A User Prompt for DeepSeek-Distilled Reasoning Models
@@
Appendix B Details of Reasoning Chain Generation
This two-stage process yields a high-quality distillation set. We deliberately choose the order—first Claude, then O3—based on qualitative observations: Claude excels at solving easier tasks and maintaining attention to safety considerations, whereas O3 performs better on harder tasks but tends to overemphasize helpfulness at the expense of safety. We select approximately 12% of the training data (slightly fewer than 9K examples) for distillation. This is then followed by RL training.
Appendix C Group Relative Policy Optimization (GRPO)
@@
$\begin{split}\mathcal{J}_{\text{GRPO}}(\theta)=\,&\mathbb{E}_{x\sim\mathcal{D},\,\{j_{i}\}_{i=1}^{G}\sim r_{\theta_{\text{old}}}(j\mid x)}\bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|j_{i}|}\sum_{t=1}^{|j_{i}|}\Big{\{}\min\Big{(}\frac{r_{\theta}(j_{i,t}\mid x,j_{i,<t})}{r_{\theta_{\text{old}}}(j_{i,t}\mid x,j_{i,<t})}\hat{A}_{i,t},\\[4.0pt] &\text{clip}\big{(}\frac{r_{\theta}(j_{i,t}\mid x,j_{i,<t})}{r_{\theta_{\text{old}}}(j_{i,t}\mid x,j_{i,<t})},1-\epsilon,1+\epsilon\big{)}\hat{A}_{i,t}\Big{)}-\beta\,\mathbb{D}_{\mathrm{KL}}\left[r_{\theta}(\cdot\mid x)\,\|\,\pi_{\text{ref}}(\cdot\mid x)\right]\Big{\}}\bigg{]},\end{split}$ (8)
where $\beta$ is a hyperparameter balancing the task specific loss and the KL-divergence. Specifically, $\hat{A}_{i}$ is computed using the rewards of a group of responses within each group $\{r_{1},r_{2},\ldots,r_{G}\}$, and is given by the following equation:
@@
Appendix D Experiment Setups
D.1 Benchmarks
@@
RewardBench [16]: RewardBench is one of the first endeavors towards benchmarking reward models through prompt-chosen-rejected trios, covering four categories: chat, chat-hard, reasoning, and safety, with 358, 456, 740, and 1431 samples, respectively.
@@
D.2 Preference Datasets
We consider the following datasets for training:
@@ D.2 Preference Datasets
A global statistics of our training dataset is summarized in Table 5.
Source Size Domain
magpile_pro_llama3.1* 29682 Reasoning
offset_bias* 8504 Chat (length bias)
helpsteer2* 7221 Chat
wildguard* 6709 Safety
magpile_pro* 2030 Chat
Code-Preference-Pairs 8000 Reasoning
Math-DPO-10K 10000 Reasoning
D.3 Baselines
@@
GenRMs. Generative reward models provide more nuanced feedback by generating textual judgments. Examples include Llama [11], Qwen [35], Claude [3], GPT-4 variants [1, 14], Gemini-1.5-pro [24], and Self-taught Evaluators [33]. GenRMs leverage generative capabilities of LLMs to produce explanations and rationales alongside numerical scores, enhancing interpretability and decision transparency.
ReasRMs. Reasoning-enhanced reward models explicitly incorporate reasoning processes into their judgments, often trained through critiques or chain-of-thought strategies. Notable examples are Critique-RM [36], DeepSeek-GRM [21], and our proposed RM-R1 models. These models excel in tasks demanding rigorous reasoning, safety evaluations, and nuanced preference judgments due to their grounding in structured critical thinking.
Appendix E Implementation Details
Our training framework is based on VERL [28] and OpenRLHF [13]. For Instruct models, we use 8.7k data for distillation and 64k for RLVR. For Deepseek-Distilled models, we use the full data for RLVR.
[…]
Appendix F Full Experiment Result
In this section, we provide the full experiment results and a more comprehensive coverage of existing baselines. The results of RewardBench, RM-Bench, and RMB are provided in Table 6Table 7Table 8, respectively.
[…]
@@ Appendix F Full Experiment Result
RM-R1-Qwen-Instruct-32B 95.3 83.1 91.9 95.2 91.4
@@
Llama3.1-70B-Instruct 64.3 67.3 47.5 83.0 74.7 67.8 54.1 65.5
Skywork-Critic-Llama-3.1-70B 71.4 64.6 56.8 94.8 85.6 73.7 56.5 71.9
GPT-4o-0806 67.2 67.5 63.6 91.7 83.4 75.6 58.7 72.5
Gemini-1.5-pro 71.6 73.9 63.7 91.3 83.1 77.6 64.7 75.2
@@
Appendix G Supplementary Information for Section 4
G.1 Ablation Settings
Cold Start RL. This approach generally involves pure RL, with rule-based rewards centered on answer correctness and format compliance. Such strategies have achieved notable success in advanced mathematical problem solving [27].
@@ G.1 Ablation Settings
$\mathcal{R}_{\text{format}}=\begin{cases}1&\text{ if format matches,}\\ 0&\text{ otherwise,}\end{cases}\quad\text{and}\quad\mathcal{R}_{\text{answer}}=\begin{cases}1&\text{ if answer matches,}\\ 0&\text{ otherwise.}\end{cases}$ (10)
@@
The total reward is the sum $\mathcal{R}=\mathcal{R}_{\text{answer}}+\mathcal{R}_{\text{format}}$. We use the prompt template shown in Figure 7, a version without any guidance on structured reasoning.